How do i write this inequality:
Your quiz grades are 19,17,20, and 15. What is the lowest grade you can receive on the next quiz and maintain at least an 18 average?
The inequality does not have to be solved. I think the answer is 71+ x /5 >=18.
Is this correct and how do i solve inequalities like this?

Answer by checkley77(12844)

71+ x /5 >=18.
x/5>=18-71
x/5>=-53
x>=-53*5
x>=-265 answer.
proof:
71-265/5>=18
71-53>=18
18>=18